OpenStudy (anonymous):

Can anyone help me set up an equation for this problem please? Traveling at full speed on a river, it takes 2 hrs for a motorboat to travel 8 miles downstream and 4 miles back upstream. If the current's speed is 2 mph, what is the maximum speed of the boat in still water?

OpenStudy (dumbcow):

Let s be boats speed. d = r*t 8 = 2(s+2) 8 = 2s +4 s = 2 however that cannot be boats max speed or else it would never be able to get upstream 4 = 2(s-2) 4 = 2s-4 s = 4 max speed = 4 mph